OpenStudy (anonymous):

What is the value of y? 4y2 + 8 = 264 ±

OpenStudy (afrodiddle):

first off get rid of 8 on both sides.

OpenStudy (afrodiddle):

264 -8

OpenStudy (anonymous):

@radar @amistre64 @Michele_Laino

OpenStudy (anonymous):

then whats> ± <

OpenStudy (afrodiddle):

O_O

OpenStudy (afrodiddle):

That's wrong.

OpenStudy (anonymous):

:( YANINA

OpenStudy (afrodiddle):

@cvliboi Yajna is wrong

OpenStudy (anonymous):

ik lol

OpenStudy (anonymous):

SHE BLOCKED ME AFTER 8 MEDALS

OpenStudy (afrodiddle):

why?

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

The question begs if you're looking for \[4y^2 +8 = 264\] or your original equation.

OpenStudy (afrodiddle):

did she give you the wrong answers?

OpenStudy (afrodiddle):

@cvliboi Subtract 8 from both sides.

OpenStudy (anonymous):

1 WRONG IN CRODENT GEMORTY

OpenStudy (anonymous):

oh sorry yea that's the equation @_Squish

OpenStudy (afrodiddle):



OpenStudy (afrodiddle):

o.o

OpenStudy (anonymous):

Here are the steps you'll like to solve in order to get your equation, as probably mentioned by previous users (I'm a very lazy person therefore I don't check all the comments above) 1. You'll want to remove 8 from both the left and right side of the equation. 2. After that, division of 256 / 4 will get you a step closer. 3. Find the square root of 64 to finalize your answer of 8. Turns out Yajna was right.

OpenStudy (afrodiddle):

@cvliboi to demonstrate exponents use "^" Such as 4y^2

OpenStudy (anonymous):

I thought it would do it by itself @Afrodiddle

OpenStudy (anonymous):

i know @_Squish

OpenStudy (afrodiddle):

nope :P

OpenStudy (afrodiddle):

Sorry @yajna Btw... why did you block that one person? o.o

OpenStudy (radar):

\[y^{2}=256/4\]\[y ^{2}=64\]\[y =\pm 8\]

OpenStudy (radar):

Verify that both plus and minus 8 satisfies the equation.

OpenStudy (er.mohd.amir):

4y^2 + 8 =264 4 y^2 =264-8 4 y^2 =256 y^2 =256/4 y*y=64 y*y=8*8 y=8 or y*y=-8*-8 y=-8 ans 8 or -8